# Muon beams towards muonium physics: progress and prospects

### v1

@@ Abstract
Advances in accelerator technology have led to significant improvements in the quality of muon beams over the past decades. Investigations of the muon and muonium enable precise measurements of fundamental constants, as well as searching for new physics beyond the Standard Model. Furthermore, by utilizing muon beams with high intensity and polarization, studies of the dynamics of the muon and muonium within atom level can offer valuable insights into material science. This review presents recent progress and prospects at the frontiers of muon beams and high-precision muonium physics. It also provides an overview of novel methods and detection techniques to achieve high sensitivities in different areas, including particle physics, nuclear physics, materials science and beyond.
@@ I Introduction
Muon was first discovered as a component of cosmic rays by C.D. Anderson and S.H. Neddermeyer in 1936 [Anderson:1936zz, Neddermeyer:1937md], and research in muon physics has remained active ever since. In the Standard Model, it is a second-generation charged lepton with a mass of 105.66 MeV [Tiesinga:2021myr], about 200 times larger than that of the electron. Muon is an unstable particle with a lifetime of about 2.2 s [MuLan:2012sih], shorter only than that of the neutron ( s) [ParticleDataGroup:2024cfk] among unstable particles. These properties make the muon an excellent probe for frontier studies in both fundamental physics and applications. Muon predominantly decays through the weak interaction into an electron and two neutrinos. For a positive muon, the normal decay mode is expressed as
@@
where the energy distribution of decay positrons follows the Michel spectrum with a maximum energy of 52.8 MeV. Due to the parity violation in weak interaction [Lee:1956qn, Wu:1957my], the angular distribution of decay positrons is asymmetrical and concentrated along the muon’s spin direction. In 1957, R.L. Garwin et al. observed parity violation in muon decay and suggested that the resulting asymmetry could serve as a powerful tool to probe magnetic fields in nuclei, atoms, and interatomic regions [Garwin:1957hc]. This work effectively predicted the wide applications of the muon spin rotation/relaxation/resonance (SR) technique nowadays [amato2024introduction, Nuccio2014, Hillier2022]. Other rare decay modes allowed by the Standard Model including and also exist, but suffer from tiny branching ratios [MEG:2016leq, SINDRUM:1985vbg]. Interestingly, some neutrinoless decay or conversion modes of muon, also known as the charged lepton flavor violation processes, are predicted by various new physics models [Calibbi:2017uvl].
@@
There are two approaches that people can obtain muons—the cosmic-ray shower and the accelerator. The dominant muon-producing channel is the two-body decay of a pion:
@@
In the case of a pion decaying at rest, the spin of is anti-parallel to its momentum direction since there are no right-handed neutrinos in the Standard Model. It enables the production of muons with a polarization close to 100% in certain conditions. Due to the interactions with the atmosphere, the cosmic-ray muon suffers from a relatively low flux (), and a highly broad energy distribution. Besides, the polarization of cosmic-ray muons also significantly decreases before they reach the sea level [sun2025cosmicraymuonpolarization]. With the development of accelerator technology, it has become possible for accelerators to reach the energy threshold for pion production since 1948 [Gardner:1948pkt]. To meet the rapidly increasing demand from the muon science community, muon beamlines based on high-intensity accelerators have been developed over the past decades. Taking advantage of the trend toward more focused energy and beam spot distribution, higher fluxes and polarization, research involving muon beams has been further extended to areas beyond particle physics [Gorringe:2015cma, Crivelli:2018vfe, Atanov:2018ich, Abe:2019thb, Mu3e:2020gyw, Accettura:2023ked, Bai:2024skk, MuSEUM:2025cmo, MEGII:2025gzr, Muong-2:2025xyk, Adelmann:2025nev], including applications in material science [ito2019investigation, nuccio2014muon, mcclelland2020muon, blundell2025muon], tomography [yu2024proposed, gamage2022portable, paccagnella2025exploring, Xu:2024btf, Yu:2025ufj, Ning:2025hai]. Evidently, the development of muon beams has served as an essential gateway to muon science.
@@
Muonium () is a unique system in muon physics [hughes2012muon]. It is a Coulombic bound state, which could be considered as an isotope of hydrogen, consisting of two point-like leptons, a positive muon and an electron . The transitions between different states can be induced by both microwave and laser spectroscopy. In 1960, V.W. Hughes et al. first observed the formation of muonium in pure argon gas via its characteristic Larmor precession frequency [Hughes:1960zz]. To date, a broad platform for muonium physics has been established. As a purely electromagnetic two-body system free from any hadronic components, it avoids the complexities associated with hadronic and nuclear structures present in ordinary atoms. Hence, the energy levels of muonium as shown as Fig. 1 can be completely calculated by Quantum Electrodynamics (QED). Compared to positronium, muonium has a higher mass and longer lifetime. Therefore, it is an ideal probe in cutting-edge research areas, including atomic spectroscopy and precision tests of quantum electrodynamics (see also comprehensive reviews of K.P. Jungmann [Jungmann:2004sa, Jungmann:2016gak]) and searches for physics beyond the Standard Model. Additionally, in the SR investigation of semiconductors or insulators, muonium is more likely to be formed. It is an appropriate system for investigating weak magnetic fields and hydrogen impurity in such materials [PhysRevB.92.081202]. In physical chemistry, muonium plays a crucial role in characterizing free radicals by SR [McKenzie2013].
@@ II.1 Overview
In most of the current and planned accelerator muon sources, muons are typically produced via pion decays, where pions are generated by the interaction of high-energy (over hundreds of MeV) protons with a target [nagamine2003introductory]. Considering a decaying at the surface of the target, the escaping is called a surface muon, which is monoenergetic (29.8 MeV/) and almost 100% polarized. By further moderating surface muons using a solid noble gas mixture [Prokscha:2008zz] or laser ionization [Nagatomo:2014bja], slow muons with energies of a few keV can be produced. Another scenario is that a pion escapes the target and decays in flight. The produced muon is called a decay muon or cloud muon, which has a higher energy (over 4.1 MeV) than the surface muon. The typical setup of a muon beamline can be described by Fig. 2. After a high-energy, high-intensity primary beam hits the production target, muons are produced with a large emittance. Therefore, it is essential to design an appropriate capture solenoid to achieve the required muon intensity. The captured muons are then transported to the experimental stations by dipole magnets for the bending purpose and quadrupole-magnet triplets for transverse focusing.
@@
In the field of muon physics, surface muon beams have found the widest range of applications. The invention of the surface muon beam dates back to the 1970s. At that time, T. Bowen et al. stood out to search for the muonium-to-antimuonium conversion process suggested by G. Feinberg and S. Weinberg [Feinberg:1961zz, Feinberg:1961zza]. To form muonium, muons have to be stopped in the production target. However, only decay muons with large momenta were available in the original beamline design. T. Bowen proposed that muons produced from pion decaying at rest in a thin target are monoenergetic and almost 100% polarized, carrying low momenta as shown in Fig. 3. Eventually, they succeeded in constructing the first surface muon beamline at Lawrence Berkeley National Laboratory (LBNL) [Pifer:1976ia, Bowen:1985qf].
@@
Accelerator facilities capable of delivering high-intensity protons or other particles are essential for the construction of muon beamlines. The major facilities currently in operation worldwide include Paul Scherrer Institute (PSI) in Switzerland, J-PARC (Japan Proton Accelerator Research Complex) in Japan, Research Center for Nuclear Physics (RCNP) in Japan, Rutherford Appleton Laboratory (RAL) in the UK, TRIUMF in Canada, and Fermi National Accelerator Laboratory (FNAL or Fermilab) in the USA, all of which are supported by advanced accelerator complexes. After years of progress in muon science, efforts are now underway to develop next-generation muon beam facilities in this century. Ongoing construction projects for new accelerator complexes include the China Spallation Neutron Source (CSNS), the China initiative Accelerator Driven System (CiADS), High Intensity heavy-ion Accelerator Facility (HIAF) in China, Rare isotope Accelerator complex for ON-line experiment (RAON) in Korea, Spallation Neutron Source (SNS) in the USA, and future muon collider concept at European Spallation Source (ESS) in Sweden [Abele:2022iml, Alekou:2022emd]. All of these facilities have proposed plans to construct dedicated muon sources, advancing the frontiers of muon science. Geographical locations and general features of these muon facilities are plotted on a world map in Fig. 4. This section provides a brief review of the key parameters and scientific cases of some major muon sources, spanning developments from the recent past to the foreseeable future.
@@ II.2 SS at PSI
The Swiss Muon Source (SS) is based on the PSI High Intensity Proton Accelerator (HIPA), which operates at a beam current of up to 2.4 mA with a power of up to 1.4 MW [Grillenberger:2021kyv]. The protons injected into the Ring cyclotron are accelerated by 50.6 MHz cavities to 590 MeV and finally sent to the meson production target stations, Target M and Target E, as plotted in Fig. 5 [Kiselev:2021pwl]. Both targets are made of polycrystalline graphite and rotate at 1 turn per second for heat dissipation. Target M measures 320 mm in diameter and 5.2 mm thick, while Target E is 450 mm in diameter and 40 mm/60 mm thick. These target stations provide pions or muons for seven secondary beamlines, including M1, M3, E1, E3, E5, E1, and E4. The detailed performances of these beamlines are listed in Table 1.
@@ II.3 MUSE at J-PARC
The J-PARC consists of a linear accelerator, a rapid-cycling synchrotron (RCS), and a main ring (MR) [Nagamiya:2012tma]. The RCS finally provides protons with an energy of 3 GeV at a 25 Hz repetition rate for the Materials and Life Science Experimental Facility (MLF), where the MUon Science Establishment (MUSE) is located. MUSE utilizes a rotating graphite target similar to the SS’s to produce muons, and contains four muon beamlines, including D-line, U-line, S-line, and H-line [Shimomura:2024puh], as plotted in Fig. 6.
The D-line delivers both positive/negative decay muons and surface muons to D1 and D2 areas. The high-energy (3 GeV) proton beam produces a high yield of negative muons. Recent applications include non-destructive elemental analysis, precision spectroscopy of muonic atoms, and related areas. Slow muons with energies ranging from 50 eV to 30 keV can be produced at the U-line through laser ionization of muonium, which is formed by stopping surface muons in materials [Kanda:2023gqp]. This opens up the opportunity for conducting research in areas similar to those pursued at the LEM facility. At the S-line, the pulsed-wave (PW) surface muon beam with a flux of is used for SR studies. Worth mentioning, a first RF acceleration of muons has been achieved at this beamline [Aritome:2024rlu]. In the other area of S-line, measurement of muonium 1S-2S transition using laser spectroscopy will also be carried out [Zhang:2021cba]. The H-line offers higher beam transport efficiency through the use of solenoids, making it suitable for the future cLFV searching experiment—Coherent Muon to Electron Transition (COMET) [COMET:2018auw], Muonium Spectroscopy Experiment Using Microwave (MuSEUM) [MuSEUM:2025cmo], muon cooling [Kamioka:2023xob], and muon measurement [Abe:2019thb].
@@ II.4 MuSIC at RCNP
The MUon Science Innovative Channel (MuSIC) facility was constructed at the RCNP, Osaka University, to demonstrate the superconducting solenoid magnets capture scheme [Hino:2014bpx, Cook:2016sfz], as shown in Fig. 7. At MuSIC, a 400 W proton cyclotron accelerator accelerates the protons to 400 MeV, which are then delivered to a graphite cylindrical target with a length of 20 cm and a radius of 2 cm. The target is located at the center of the capture solenoid (PCS).
@@ II.5 ISIS at RAL
The ISIS Pulsed Neutron and Muon Source, located at the RAL, has operated muon facilities for over 30 years [Thomason:2019fwe]. The ISIS proton accelerator consists of a Radio Frequency Quadrupole (RFQ), a linear accelerator, and a synchrotron, delivering 800 MeV protons to the muon target station at 50 Hz. The target is made of graphite and has a thickness of 10 mm along the beam direction. The muon target station provides 28 MeV/ surface muons to the European Commission facility, as well as decay muons (positive or negative) with momenta ranging from 17 to 120 MeV/ to the RIKEN-RAL facility [Matsuzaki:2001kj, Hillier:2019szm]. There is also a dedicated port for slow muon research at the RIKEN-RAL facility, but the intensity is rather small [Matsuda:2001vhp]. The layout of the instruments at ISIS Pulsed Neutron and Muon Source is shown as Fig. 8.
@@ II.6 CMMS at TRIUMF
The first extraction of the beam from the TRIUMF cyclotron took place in 1974. It delivers 520 MeV protons to the muon production targets T1 and T2 with a beam power of 75 kW [beveridge1986muon, 987839] to the Centre for Molecular and Materials Science (CMMS). Detailed beamline layout can be seen in Fig. 9. The M15 beamline provides surface muons from the 1AT1 target, typically made of 1 cm graphite or beryllium. A permanent quadrupole doublet located close to the target enables efficient muon collection, but limits the beam momentum range to approximately 19–40 MeV/. M20 extracts surface muons from a 10 cm thick beryllium block target called 1AT2. The M9 beamline branches from the 1AT2 target into two legs. M9A, equipped with an ultra-fast electrostatic, is being rebuilt as a surface muon beamline dedicated to SR experiments. M9H will set up a persistent superconducting solenoid where pions decay in flight to produce high-energy and beams. It is optimized for transverse-field SR measurements across a wide energy range. It will also support high-pressure liquid or gas sample environments under extreme conditions (1000 K and 0.6 GPa).

### v2

@@ Abstract
Advances in accelerator technology have led to significant improvements in the quality of muon beams over the past decades. Investigations of the muon and muonium enable precise measurements of fundamental constants, as well as searches for new physics beyond the Standard Model. Furthermore, by utilizing muon beams with high intensity and polarization, studies of the dynamics of the muon and muonium within atomic level can offer valuable insights into materials science. This review presents recent progress and prospects at the frontiers of muon beams and high-precision muonium physics. It also provides an overview of novel methods and detection techniques to achieve high sensitivities in different areas, including particle physics, nuclear physics, materials science and beyond.
Contents
@@
- II Status of muon beams
- III Future prospects of muon beams
- IV Experimental studies on muonium physics
- V Applications of muon beams
- VI Summary and outlook
I Introduction
@@
Muon was first observed as a component of cosmic rays by C.D. Anderson and S.H. Neddermeyer in 1936 [Anderson:1936zz, Neddermeyer:1937md], and research in muon physics has remained active ever since. In the Standard Model, it is the second-generation charged lepton with a mass of 105.66 MeV [Tiesinga:2021myr], about 200 times larger than the electron. Muon is an unstable particle with a lifetime of about 2.2 s [MuLan:2012sih], shorter only than that of the neutron ( s) [ParticleDataGroup:2024cfk]. These properties make the muon an excellent probe for frontier studies in both fundamental physics and applications. Muon predominantly decays through the weak interaction into an electron and two neutrinos. For the muon, the normal decay modes are expressed as
@@
where the energy distribution of decay positrons follows the Michel spectrum with a maximum energy of 52.8 MeV. Due to the parity violation in weak interaction [Lee:1956qn, Wu:1957my], the angular distribution of decay positrons is asymmetrical and concentrated along the muon’s spin direction. In 1957, R.L. Garwin et al. observed parity violation in muon decay and suggested that the resulting asymmetry could serve as a powerful tool to probe magnetic fields in nuclei, atoms, and interatomic regions [Garwin:1957hc]. This work effectively predicted the wide applications of the muon spin rotation/relaxation/resonance (SR) technique nowadays [amato2024introduction, Nuccio2014, Hillier2022]. Other rare decay modes allowed by the Standard Model, which include and also exist, but suffer from tiny branching ratios [MEG:2016leq, SINDRUM:1985vbg]. Interestingly, some neutrinoless decay or conversion modes of muon, also known as the charged lepton flavor violation processes, are predicted by various new physics models [Calibbi:2017uvl].
@@
There are two approaches by which one can obtain muons—the cosmic-ray showers and the accelerators. The dominant muon-producing channel is the two-body decay of a pion:
@@
In the case of a pion decaying at rest, the spin of is anti-parallel to its momentum direction since there are no right-handed neutrinos in the Standard Model. It enables the production of muons with a polarization close to 100% in certain conditions. Due to the interactions with the atmosphere, the cosmic-ray muon suffers from a relatively low flux (), and a highly broad energy distribution. Besides, the polarization of cosmic-ray muons also significantly decreases before they reach the sea level [sun2025cosmicraymuonpolarization]. With the development of accelerator technology, it has become possible for accelerators to reach the energy threshold for pion production since 1948 [Gardner:1948pkt]. To meet the rapidly increasing demand from the muon science community, muon beamlines based on high-intensity accelerators have been developed over the past decades. Taking advantage of the trend toward more focused energy and beam spot distribution, higher fluxes and polarization, research involving muon beams has been further expanded to areas beyond particle physics [Crivelli:2018vfe, Atanov:2018ich, Abe:2019thb, Mu3e:2020gyw, Accettura:2023ked, Bai:2024skk, MuSEUM:2025cmo, MEGII:2025gzr, Muong-2:2025xyk, Adelmann:2025nev], including applications in materials science [ito2019investigation, nuccio2014muon, mcclelland2020muon, blundell2025muon], tomography [yu2024proposed, gamage2022portable, paccagnella2025exploring, Xu:2024btf, Yu:2025ufj, Ning:2025hai]. Evidently, the development of muon beams has served as an essential gateway to muon science.
@@
Muonium () is a unique system in muon physics [hughes2012muon]. It is a Coulombic bound state, which could be considered as an isotope of hydrogen, consisting of two point-like leptons, a positive muon and an electron . The transitions between different states can be induced by both microwave and laser spectroscopy. In 1960, V.W. Hughes et al. first observed the formation of muonium in pure argon gas via its characteristic Larmor precession frequency [Hughes:1960zz]. To date, a broad platform for muonium physics has been established. As a purely electromagnetic two-body system free from any hadronic components, it avoids the complexities associated with hadronic and nuclear structures present in ordinary atoms. Hence, the energy levels of muonium as shown in Fig.˜1 can be completely calculated by Quantum Electrodynamics (QED). Compared to positronium, muonium has a higher mass and a longer lifetime. Therefore, it is an ideal probe in cutting-edge research areas, including atomic spectroscopy, precision tests of quantum electrodynamics and searches for physics beyond the Standard Model (see also comprehensive reviews of K.P. Jungmann [Jungmann:2004sa, Jungmann:2016gak] and T.P. Gorringe [Gorringe:2015cma]). Additionally, in the SR investigation of semiconductors or insulators, muonium is more likely to be formed. It is an appropriate system for investigating weak magnetic fields and hydrogen impurity in such materials [PhysRevB.92.081202]. In physical chemistry, muonium plays a crucial role in characterizing free radicals by SR [McKenzie2013].
@@ II.1 Overview
In most of the current and planned accelerator muon sources, muons are typically produced via pion decays, where pions are generated by the interaction of high-energy (over hundreds of MeV) protons with a target [nagamine2003introductory]. Considering a decaying at the surface of the target, the escaping is called a surface muon, which is monoenergetic (29.8 MeV/) and almost 100% polarized. By further moderating surface muons using a solid noble gas mixture [Prokscha:2008zz] or laser ionization [Nagatomo:2014bja], slow muons with energies of a few keV can be produced. Another scenario is that a pion escapes the target and decays in flight. The produced muon is called a decay muon or cloud muon, which has a higher energy (over 4.1 MeV) than the surface muon. The typical setup of a muon beamline can be described by Fig.˜2. After a high-energy, high-intensity primary beam hits the production target, muons are produced with a large emittance. Therefore, it is essential to design an appropriate capture solenoid to achieve the required muon intensity. The captured muons are then transported to the experimental stations by dipole magnets for the bending purpose and quadrupole-magnet triplets for transverse focusing. According to the time structure of the primary beam, muon beams can be classified as continuous-wave (CW) or pulsed-wave (PW).
@@
In the field of muon physics, surface muon beams have found the widest range of applications. The invention of the surface muon beam dates back to the 1970s. At that time, T. Bowen et al. stood out to search for the muonium-to-antimuonium conversion process suggested by G. Feinberg and S. Weinberg [Feinberg:1961zz, Feinberg:1961zza]. To form muonium, muons have to be stopped in the production target. However, only decay muons with large momenta were available in the original beamline design. T. Bowen proposed that muons produced from pions decaying at rest in a thin target are monoenergetic and almost 100% polarized, carrying low momenta as shown in Fig.˜3. Eventually, they succeeded in constructing the first surface muon beamline at Lawrence Berkeley National Laboratory (LBNL) [Pifer:1976ia, Bowen:1985qf].
@@
Accelerator facilities capable of delivering high-intensity protons or other particles are essential for the construction of muon beamlines. The major facilities currently in operation worldwide include Paul Scherrer Institute (PSI) in Switzerland, J-PARC (Japan Proton Accelerator Research Complex) in Japan, Research Center for Nuclear Physics (RCNP) in Japan, Rutherford Appleton Laboratory (RAL) in the UK, TRIUMF in Canada, and Fermi National Accelerator Laboratory (FNAL or Fermilab) in the USA, all of which are supported by advanced accelerator complexes. After years of progress in muon science, efforts are now underway to develop next-generation muon beam facilities in this century. Ongoing construction projects for new accelerator complexes include the China Spallation Neutron Source (CSNS), the China initiative Accelerator Driven System (CiADS), High Intensity heavy-ion Accelerator Facility (HIAF) in China, Rare isotope Accelerator complex for ON-line experiment (RAON) in Korea, Spallation Neutron Source (SNS) in the USA, and future muon collider concept at European Spallation Source (ESS) in Sweden [Abele:2022iml, Alekou:2022emd]. All of these facilities have proposed plans to construct dedicated muon sources, advancing the frontiers of muon science. Geographical locations and general features of these muon facilities are plotted on a world map in Fig.˜4. This section provides a brief review of the key parameters and scientific cases of some major muon sources, spanning developments from the recent past to the foreseeable future.
@@ II.2 SS at PSI
The Swiss Muon Source (SS) is based on the PSI High Intensity Proton Accelerator (HIPA), which operates at a beam current of up to 2.4 mA with a power of up to 1.4 MW [Grillenberger:2021kyv]. The protons injected into the Ring cyclotron are accelerated by 50.6 MHz cavities to 590 MeV and finally sent to the meson production target stations, Target M and Target E, as plotted in Fig.˜5 [Kiselev:2021pwl]. Both targets are made of polycrystalline graphite and rotate at 1 turn per second for heat dissipation. Target M measures 320 mm in diameter and 5.2 mm thick, while Target E is 450 mm in diameter and 40 mm or 60 mm thick. These target stations provide pions or muons for seven secondary beamlines, including M1, M3, E1, E3, E5, E1, and E4. The detailed performances of these beamlines are listed in Table˜1.
@@ II.3 MUSE at J-PARC
The J-PARC consists of a linear accelerator, a rapid-cycling synchrotron (RCS), and a main ring (MR) [Nagamiya:2012tma]. The RCS finally provides protons with an energy of 3 GeV at a 25 Hz repetition rate for the Materials and Life Science Experimental Facility (MLF), where the MUon Science Establishment (MUSE) is located. MUSE utilizes a rotating graphite target similar to the SS’s to produce muons, and contains four muon beamlines, including D-line, U-line, S-line, and H-line [Shimomura:2024puh], as plotted in Fig.˜6.
The D-line delivers both positive/negative decay muons and surface muons to D1 and D2 areas. The high-energy (3 GeV) proton beam produces a high yield of negative muons. Recent applications include non-destructive elemental analysis, precision spectroscopy of muonic atoms, and related areas. Slow muons with energies ranging from 50 eV to 30 keV can be produced at the U-line through laser ionization of muonium, which is formed by stopping surface muons in materials [Kanda:2023gqp]. This opens up the opportunity for conducting research in areas similar to those pursued at the LEM facility. At the S-line, the PW surface muon beam with a flux of is used for SR studies. It is worth mentioning that the first RF acceleration of muons has been achieved at this beamline [Aritome:2024rlu]. In the other area of S-line, the measurement of muonium 1S-2S transition using laser spectroscopy will also be carried out [Zhang:2021cba]. The H-line offers higher beam transport efficiency through the use of solenoids, making it suitable for the future cLFV search experiment—Direct Emission of Electrons from Muon to Electron conversions (DeeMe) [Natori:2014yba], Muonium Spectroscopy Experiment Using Microwave (MuSEUM) [MuSEUM:2025cmo], muon cooling and acceleration [Kamioka:2023xob], muon measurement [Abe:2019thb], and the world’s first muon accelerator [Aritome:2024rlu].
@@ II.4 MuSIC at RCNP
The MUon Science Innovative Channel (MuSIC) facility was constructed at the RCNP, the University of Osaka, to demonstrate the superconducting solenoid magnets capture scheme [Hino:2014bpx, Cook:2016sfz], as shown in Fig.˜7. At MuSIC, a 400 W proton cyclotron accelerator accelerates the protons to 400 MeV, which are then delivered to a graphite cylindrical target with a length of 20 cm and a radius of 2 cm. The target is located at the center of the capture solenoid (PCS).
@@ II.5 ISIS at RAL
The ISIS Pulsed Neutron and Muon Source, located at the RAL, has operated muon facilities for over 30 years [Thomason:2019fwe]. The proton accelerator consists of a Radio Frequency Quadrupole (RFQ), a linear accelerator, and a synchrotron, delivering 800 MeV protons to the muon target station at 50 Hz. The target is made of graphite and has a thickness of 10 mm along the beam direction. The muon target station provides 28 MeV/ surface muons to the European Commission facility, as well as decay muons (positive or negative) with momenta ranging from 17 to 120 MeV/ to the RIKEN-RAL facility [Matsuzaki:2001kj, Hillier:2019szm]. There is also a dedicated port for slow muon research at the RIKEN-RAL facility, but the intensity is rather low [Matsuda:2001vhp]. The layout of the instruments at ISIS Pulsed Neutron and Muon Source is shown in Fig.˜8.
@@ II.6 CMMS at TRIUMF
The first extraction of the beam from the TRIUMF cyclotron took place in 1974. It delivers 520 MeV protons to the muon production targets T1 and T2 with a beam power of 75 kW [beveridge1986muon, 987839] to the Centre for Molecular and Materials Science (CMMS). Detailed beamline layout can be seen in Fig.˜9. The M15 beamline provides surface muons from the 1AT1 target, typically made of 1 cm graphite or beryllium. A permanent quadrupole doublet located close to the target enables efficient muon collection, but limits the beam momentum range to approximately 19–40 MeV/. M20 extracts surface muons from a 10 cm thick beryllium block target called 1AT2. The M9 beamline branches from the 1AT2 target into two legs. M9A, equipped with an ultra-fast electrostatic, is being rebuilt as a surface muon beamline dedicated to SR experiments. M9H will set up a persistent superconducting solenoid where pions decay in flight to produce high-energy and beams. It is optimized for transverse-field SR measurements across a wide energy range. It will also support high-pressure liquid or gas sample environments under extreme conditions (1000 K and 0.6 GPa).